Mary is 24. she is twice as old as Ana was when Mary as old as Ana is now. How old is Ana?
Answer by robertb(5830)

Let a = Ana's age now.The difference between their ages NOW is 24-a.
24 - a years ago, Mary was a years old, while Ana was a-(24-a) = 2a - 24.
During that time, a = 2(2a-24), or a = 4a - 48, or 3a = 48, or a = 16.
Thus Ana is 16 years old now.